Cassini Significant Events 02/03/05 - 02/09/05

February 11, 2005

(Source: Jet Propulsion Laboratory)


Cassini Significant Events
for 02/03/05 - 02/09/05


The most recent spacecraft telemetry was acquired today from the Goldstone
tracking station. The Cassini spacecraft is in an excellent state of health
and is operating normally. Information on the present position and speed of
the Cassini spacecraft may be found on the "Present Position" web page
located at http://saturn.jpl.nasa.gov/operations/present-position.cfm .


Activities this week:


We are now in the third week of execution for tour sequence S08. The
Imaging Science Subsystem (ISS) and the Visible and Infrared Mapping
Spectrometer (VIMS) conducted observations for time variable atmospheric
events as we approach Saturn. In this last week 797 ISS images arrived
along with 127 VIMS cubes.


Each day this week the Optical Remote Sensing instruments are taking data,
the Navigation team is obtaining Optical Navigation images, and we have - on
average - one DSN pass a day where we downlink the data. If you don't see
anything listed on a particular day, it's because we are just doing business
as usual and have no extra events going on.


Thursday, February 3:


We have officially vacated the Mission Support Area on the first floor.
Cassini is staffed in the Space Flight Operations Facility at JPL. When
missions launch or arrive they may get space in the operations area on the
first floor. We were there back in 1997 when we launched, and then moved
back in for Saturn Orbit Insertion in mid 2004. Now that Probe activities
have concluded, we have once again vacated the area and the Deep Impact
Project is expected to move in shortly in preparation for their upcoming
encounter. Cassini teams are now supporting commanding from their local
office areas.


On-board the spacecraft today we ran a periodic engineering maintenance
activity, and an Ion and Neutral Mass Spectrometer planned reboot. There
is a second reboot scheduled for Sunday of this week as well.


On the ground, the Preliminary Sequence Integration and Validation 2 merge
products for S09 were released for review, and the Ground Software Monthly
Management Review was held. This is where the teams get together and report
status on software in development, and future plans for deliveries.
Currently the primary delivery will occur in May of this year.


Friday, February 4:


A couple of events occurred on board the spacecraft today. First, the
Spacecraft Operations Office (SCO) executed a procedure to remove the Probe
Relay sequence from the SSRs. This is the final activity in support of
Probe operations. This activity began yesterday and completed today over
Madrid's DSS #63 pass.


The second activity was to fire pyro valve #26. Firing this valve isolates
the oxidizer system and prevents any possibility of mixing of propellant
components. This is actually a clean up activity from the periapsis raise
maneuver that occurred in August 2004. Since it was not a critical
activity it was decided to delay it until after Probe Relay in order to
maintain the spacecraft in as quiet a state as possible.


A member of the Project Science staff supported an American Association of
University Women career day at Whittier College today. Presentations on
Cassini science results were given to 8th grade girls to encourage them to
continue to take math and science classes in high school.


All teams and offices supported the Cassini / NASA Quarterly review today.


Monday, February 7:


The Radio Science Subsystem (RSS) conducted an Operations Readiness Test of
their instrument today. Additional activities occur Tuesday, Thursday, and
Sunday.


A preliminary estimate of the wind variations on Titan with altitude from
about 100 km down to the surface has been recovered by a joint team of
researchers from NASA's Jet Propulsion Laboratory collaborating with the
Huygens Doppler wind experiment team, and with the ground-based European
Very Long Baseline Interferometry team. For more information check out the
Cassini web site.


Tuesday, February 8:


RSS performed an Ultra Stable Oscillator characterization and a periodic
instrument maintenance today.


SCO conducted the first of a three-part gyro calibration today. Parts B and
C will execute on Wednesday and a status report will be released next week.


Uplink Operations sent up commands to the spacecraft to make the Cosmic Dust
Analyzer the prime instrument on day of year 47 -February 16- during the
ring plane crossing. They also sent up Instrument Expanded Block files for
the Magnetospheric Imaging Instrument, and a Radar Enceladus mini-sequence.


Wednesday, February 9:


A wrap up meeting was held today for the S39/S40 Science Operations Plan
implementation process. These sequences will now be archived. They will
next be dusted off for further processing in September of 2007!


ISS presented recent science results of Iapetus at a Tour Science Talk, and
the Satellite Orbiter Science Team hosted an Enceladus preview meeting where
science objectives and activities were reviewed. The flyby is on February
17. The next Tour Science Talk is on February 16 and will be the Composite
InfraRed Spectrometer team presenting some of their Iapetus results.


Wrap up:


That's it for this week.


The outreach folks especially asked me to put in a plug for local events for
the Saturn Observation Campaign. For views of Saturn and the moon near JPL:


* Friday, February 18, Colorado Blvd. near Delacey, Pasadena

* Saturday, February 19, Myrtle & Lime, Monrovia

* Friday, March 18, Pasadena

* Saturday, March 19, Monrovia